A sample of gas has an initial volume of 27.7L at a pressure of 1.4 atm

If a sample is compressed to a volume of 11.8L what will its pressure be? Assume constant temperature.
1
Expert's answer
2017-11-13T08:54:07-0500
P1 = 1.4 atm
V1 = 27.7 l
V2 = 11.8 l
P2 - ?
P1V1 = P2V2
P2 = P1V1 / V2
P2 = 1.4*27.7/11.8 = 3.29 (atm)